The Organizing Committee of the 15-th ILAS Conference acknowledges the active participation of

Pedro Aceves

Instituto de Matemáticas Unidad Morelia, Morelia, México in all the conference activities.

Richard Akinola

University of Bath, Bath, United Kingdom in all the conference activities.

Abdulhadi Aminu

University of Birmingham, Birmingham, United Kingdom in all the conference activities.

José Antonio Arciniega-Nevarez

Instituto de Matemáticas, UNAM, Cuernavaca

in all the conference activities.

Luis Verde-Star

Ma. de Lourdes Ariza

Instituto de Matemáticas, UNAM, Cuernavaca in all the conference activities.

María José Arroyo

UAM-Iztapalapa, Mexico, City

in all the conference activities.

Luis Verde-Star

Vanesa Ávalos Gaytán

Universidad Autónoma de Nuevo León, Nuevo León, México

in all the conference activities.

Luis Verde-Star

Julio César Ávila-Romero

Instituto de Matemáticas, UNAM, Cuernavaca in all the conference activities.

Fernando Bernal-Vilchis

Instituto de Matemáticas, UNAM, Morelia in all the conference activities.

Barbara Bianchini

Pontificia Universidade Catolica de Sao Paulo, Sao Paulo in all the conference activities.

James Bunch

UCSD, La Jolla, CA in all the conference activities.

Alejandro Waldemar Coba-Magaña

Instituto de Matemáticas, UNAM, Morelia

in all the conference activities.

Luis Verde-Star

Roberto Canogar

Universidad Nacional de Educación a Distancia, Madrid, Spain

in all the conference activities.

Luis Verde-Star

Erick Coronado-Juárez

Instituto de Matemáticas, UNAM, Mexico City in all the conference activities.

Claudia Corriveau

Université du Québec à Montréal (UQAM), Montréal, Canada in all the conference activities.

Javier de Jesús Cortés-Aguirre

Facultad de Ciencias, UNAM, Mexico City

in all the conference activities.

Luis Verde-Star

Ljiljana Cvetkovic

Faculty of Science, Novi Sad, Serbia

in all the conference activities.

Luis Verde-Star

Adriana Dávila

Instituto de Matemáticas, UNAM, Mexico City in all the conference activities.

Biswajit Deb

Sikkim Manipal Institute of Technology, Majitar, Sikkim, India

in all the conference activities.

Luis Verde-Star

Alejandro Díaz-Gutiérrez

Instituto de Matemáticas, UNAM, Cuernavaca

in all the conference activities.

Luis Verde-Star

Ricardo Díaz-Gutiérrez

Instituto de Matemáticas, UNAM, Cuernavaca in all the conference activities.

Mirko Dobovisek

University of Ljubljana, Faculty of Mathematics & Physics, Ljubljana,

Slovenia in all the conference activities.

Gernot Engel

Transversal Networks Inc., Madison Wisconsin

in all the conference activities.

Luis Verde-Star

María Esperanza Guzmán Ovando

Benemerita Universidad Autónoma de Puebla, Puebla

in all the conference activities.

Luis Verde-Star

Laurent Hardouin

LISA - University of Angers, Angers, France in all the conference activities.

Magdalena Hernández-Cedillo

Instituto de Matemáticas, UNAM, Mexico City

in all the conference activities.

Luis Verde-Star

Esteban Librado Hernández Escamilla

UNAM, Mexico, D.F. in all the conference activities.

Crispín Herrera

ESFM-IPN, Mexico City

in all the conference activities.

Luis Verde-Star

Ronald Richard Jiménez-Munguía

Instituto de Matemáticas, UNAM, Cuernavaca

in all the conference activities.

Luis Verde-Star

Jorge Otero

BUAP, Puebla

in all the conference activities.

Luis Verde-Star

César Carreón-Otañez

Facultad de Ciencias, UNAM, Mexico City in all the conference activities.

Gabriel Kantún Montiel

Benemérita Universidad Autónoma de Puebla, Puebla, México in all the conference activities.

Vladimir Kostici

Faculty of Science, Novi Sad, Serbia in all the conference activities.

Darly KÃ^{o}

CINVESTAV, Distrito Federal

in all the conference activities.

Luis Verde-Star

Sangho Kum

Chungbuk National University, Cheongju, S. Korea

in all the conference activities.

Luis Verde-Star

Christine Larson

San Diego State University & Indiana University, San Diego in all the conference activities.

David Lay

University of Maryland, College Park, Maryland

in all the conference activities.

Luis Verde-Star

Yongdo Lim

Kyungpook National University, Daegu, Korea

in all the conference activities.

Luis Verde-Star

Raphael Loewy

Technion-Israel Institute of Technology, Haifa, Israel

in all the conference activities.

Luis Verde-Star

Gloria López-Chávez

Instituto de Matemáticas, UNAM, Mexico City

in all the conference activities.

Luis Verde-Star

Alma Delia López-Trujillo

Instituto de Matemáticas, UNAM, Cuernavaca

in all the conference activities.

Luis Verde-Star

Johana Luviano Flores

Instituto de Matemáticas, UNAM, Mexico City in all the conference activities.

Maria Cristina Maranhao

Pontificia Universidade Católica de SÃ£o Paulo, SÃ£o Paulo in all the conference activities.

Rubén Martínez-Avendaño

Universidad Autónoma del Estado de Hidalgo, Pachuca in all the conference activities.

Gerardo Martínez

Escuela de Ciencias UABJO, Oaxaca in all the conference activities.

Habel Mathewkutty

Habel Math, Humble, TX

in all the conference activities.

Luis Verde-Star

Russ Merris

Hayward, CA in all the conference activities.

Daniel Miguel-Ortiz

ESFM-IPN, Mexico City

in all the conference activities.

Luis Verde-Star

Pramod Kumar Mishra

BHU Varanasi ( India), Varanasi in all the conference activities.

Hossein Moosaei

Guilan, Rasht, Iran in all the conference activities.

Joel Nava-Lara

ESFM-IPN, Mexico City in all the conference activities.

Max Neumann-Coto

Instituto de Matemáticas, UNAM, Mexico City in all the conference activities.

Diego Pérez

UNAM, Mexico

in all the conference activities.

Luis Verde-Star

Reyna María Pérez-Tiscareño

Instituto de Matemáticas, UNAM, Mexico City

in all the conference activities.

Luis Verde-Star

Kernel Enrique Prieto-Moreno

IIMAS UNAM, Mexico City

in all the conference activities.

Luis Verde-Star

Luis Manuel Rivera Martínez

IICO, Universidad Autónoma de San Luis Potosí, San Luis Potosí in all the conference activities.

Solange Roa Fuentes

CINVESTAV, México Distrito Federal

in all the conference activities.

Luis Verde-Star

Robert Ross

Elsevier, Cambridge, MA in all the conference activities.

Antonio Jesús Sánchez

ESFM-IPN, Mexico City in all the conference activities.

Pedro A. Santos

Instituto Superior Técnico, UTL, Lisbon, Instituto Superior Técnico, UTL, Lisbon, Portugal

in all the conference activities.

Luis Verde-Star

Eliseo Sarmiento-Rosales

ESFM-IPN, Mexico City in all the conference activities.

Pappur Shivakumar

University of Manitoba, Winnipeg, Canada

in all the conference activities.

Luis Verde-Star

Martha Takane-Imay

Instituto de Matemáticas UNAM (Cuernavaca), Cuernavaca

in all the conference activities.

Luis Verde-Star

Kin Po Tam

University of Birmingham, Birmingham, United Kingdom in all the conference activities.

Francisco Treviño

Universidad Autónoma Metropolitana Azcapotzalco, México in all the conference activities.

Luis Carlos Velázquez-Guerrero

Facultad de Ciencias, UNAM, Mexico City

in all the conference activities.

Luis Verde-Star

Armando Esteban Villalobos-Sánchez

Facultad de Ciencias, UNAM, Mexico City in all the conference activities.